### v2

# Domain Adaptation of Foundation LLMs for e-Commerce

Abstract
We present the e-Llama models: 8 billion and 70 billion parameter large language models that are adapted towards the e-commerce domain. These models are meant as foundation models with deep knowledge about e-commerce, that form a base for instruction- and fine-tuning. The e-Llama models are obtained by continuously pretraining the Llama 3.1 base models on 1 trillion tokens of domain-specific data.
We discuss our approach and motivate our choice of hyperparameters with a series of ablation studies. To quantify how well the models have been adapted to the e-commerce domain, we define and implement a set of multilingual, e-commerce specific evaluation tasks.
We show that, when carefully choosing the training setup, the Llama 3.1 models can be adapted towards the new domain without sacrificing significant performance on general domain tasks. We also explore the possibility of merging the adapted model and the base model for a better control of the performance trade-off between domains.
Domain Adaptation of Foundation LLMs for e-Commerce
Christian Herold  Michael Kozielski  Tala Bazazo Pavel PetrushkovSeyyed Hadi HashemiPatrycja Cieplicka Dominika Basajthanks: work done while at eBayShahram Khadivi eBay Inc. {cherold, skhadivi}@ebay.com
1 Introduction
Large Language Models (LLMs) have greatly improved the performance on countless natural language processing applications across many domains Brown et al. (2020). While there exist closed services such as OpenAI’s ChatGPT OpenAI (2023), Anthropic’s Claude Anthropic (2024) and Google’s Gemini Anil et al. (2023), these are very expensive to use, don’t scale well to large amounts of requests, and pose a risk e.g. in terms of data security.
To mitigate these risks, there exists the possibility of developing LLMs from scratch to better fit the specific requirements in terms of scalability, domain knowledge, inference costs, etc. However, training competitive LLMs from scratch requires huge amounts of resources, including both compute and manpower, making it practically intractable for most teams. For example, to develop the most recent Llama-3 models at meta, more than 500 people contributed to the project and 16,000 H100 GPUs were utilized Dubey et al. (2024).
Instead of training from scratch, one could use existing pretrained models, such as Llama-3.1 Dubey et al. (2024) for their use-cases. However, these models typically lack specific knowledge, in our case about the e-commerce domain.
As a solution, we continue training the Llama base models on a large amount of e-commerce data. This way we introduce the domain specific knowledge into the model while at the same time keeping the general capabilities of the model intact. This technique is known as ‘continued pretraining‘ and the training setup has to be carefully balanced to prevent the model from degrading too much in performance on general domain tasks.
In Table 1, we compare recent continuous pretraining works in terms of the domain, the size of the models as well as the amount of training data.
Study Domain Model Parameter Count Total num Tokens
Minerva Lewkowycz et al. (2022) STEM 8B, 62B, 540B 26B-38.5B
MediTron Chen et al. (2023) Medicine 7B, 70B 46.7B
Code Llama Rozière et al. (2023) Code 7B, 13B, 34B, 70B 520B-1, [^0] B
Llemma Azerbayev et al. (2024) Math 7B, 34B 50B-55B
DeepSeekMath Shao et al. (2024) Math 7B 500B
SaulLM-7B Colombo et al. (2024b) Law 7B 30B
SaulLM-54, 141B Colombo et al. (2024a) Law 54B, 141B 520B
HEAL Yuan et al. (2024) Medicine 13B 14.9B
Me-LLaMA Xie et al. (2024) Medicine 13B, 70B 129B
ClimateGPT Thulke et al. (2024) Climate 7B, 13B, 70B 4.2B
Nemotron Parmar et al. (2024) General 15B 1, [^0] B
e-Llama (ours) e-commerce 8B, 70B 1, [^0] B
Table 1: Comparing the scale of recent continued pretraining works with our setting. Most existing works are at a significantly smaller scale, either in terms of model size or in terms of tokens used for training.
As can be seen, our work is at a significantly larger scale than most existing works, either in terms of model size or in terms of tokens used for training, or both. We share our insights regarding large-scale model adaptation. In particular, we compare the model adaptation for models of different sizes and discuss the observed differences in behavior. We also explore the possibility of model merging to better control the trade-off between general- and domain-specific knowledge.
The rest of the paper is structured as follows. In Section 2 we discuss our data mixture and explain our methods of model evaluation with focus on e-commerce specific tasks. In Section 3 we explain our series of experiments to determine the optimal set of hyperparameters. In Section 4 we show the performance of the final models and discuss the possibility of model merging to better tune for different domains.
2 Setup, Data and Evaluation
In this section, we discuss the experimental setup, with a focus on the new e-commerce benchmarks that we have defined.
2.1 Training Framework and Hardware
For training, we use the Megatron-LM framework from NVIDIA Shoeybi et al. (2019); Narayanan et al. (2021). Training was conducted using 60 nodes, each having 8 NVIDIA H100 80GB GPUs (a total of 480 GPUs). The GPUs are connected via NVIDIA NVLink (intra-node) and InfiniBand (inter-node). The hardware is part of the eBay compute platform.
2.2 Data
Regarding training data, we mostly follow Herold et al. (2024). For general domain data we use a mixture of web-crawled and smaller but more high quality datasets. We include 10% non-English general domain data in the data mix. Regarding the e-commerce domain, we utilize several data sources. On the one hand, we utilize listings and product reviews from the eBay website, as has been done by Herold et al. (2024). Furthermore, inspired by Lozhkov et al. (2024), we train an e-commerce classifier and use it to extract e-commerce specific examples from the Fineweb corpus Penedo et al. (2024). We use 20% of this data in our e-commerce specific data mixture.
2.3 Evaluation
We perform evaluation both on general and e-commerce specific tasks. As a first benchmark, we calculate model perplexity on heldout datasets for general and e-commerce data.
General Domain
For evaluating the model capabilities on the general domain for the English language, we utilize the Natural Language Understanding (NLU) benchmark aggregates (in the following called NLU En) also used by Groeneveld et al. (2024) and Herold et al. (2024) and calculated using the EleutherAI LM Evaluation Harness Gao et al. (2023). Furthermore, we utilize the ‘Open LLM Leaderboard 2’ Fourrier et al. (2024) (in the following called LLM Leaderboard) which calculates a re-normalized average of the scores for the BBH Suzgun et al. (2022), GPQA Rein et al. (2023), MUSR Sprague et al. (2024) and MMLU-PRO Wang et al. (2024) benchmarks. [^1] We exclude IFEval and MATH Lvl 5 benchmarks because the former is only useful for instruction-tuned models and the latter gives very low scores for the base models, especially for the 8B model variants. When we report model performance in Section 3, we average NLU En and LLM Leaderboard scores. For the evaluation of the non-English, general domain NLU capabilities we use the same task aggregates as Herold et al. (2024) (in the following called NLU non-En). In this work we focus on German, Spanish, French and Italian only.
e-commerce
Since existing work, like eCeLLM Peng et al. (2024) and EcomGPT Li et al. (2024) focuses on evaluation of instruction tuned models, we define a total of 5 novel e-commerce benchmarks for evaluation of foundation models. All tasks are strongly connected to relevant downstream tasks that we encounter in the e-commerce setting. They revolve around the listings on an e-commerce website, of which we consider title, category, price and a list of aspect key-value pairs [^2] An example for an aspect key could be ‘Brand’ and a possible aspect-value in this case could be ‘Nike’.. Below we list the tasks in detail:
  1. 1.
    Aspect Prediction (AP): Given the title and category of a listing, as well as a specific aspect key, predict the corresponding aspect value.
  2. 2.
[…]
  3. 3.
[…]
  4. 4.
    Most Common Aspects (MCA): Given a category and an aspect key, the model has to predict the most common aspect values for that key.
  5. 5.
    Most Common Aspects Multiple Choice (${}^{\text{MC}}$): Given a category and an aspect key, the model is presented with 4 choices for the most common aspect value for that key and has to select the correct one.
We evaluate these tasks for English, German, Spanish, French and Italian. For all tasks, the final evaluation metric is accuracy. We give an example for each of the task in Appendix A.1.
In order to obtain a strong baseline, we perform a set of experiments where we optimize the number of few-shot examples for the base Llama-3 model, see Appendix A.2 for the details.
3 Finding the best Setup
In this section we discuss several series of experiments we performed to determine the best setup for continuously pretraining. For these studies we focus on the English language benchmarks, since we assume the non-English languages will follow the same trend. Since the 3.1 version of Llama was not released at the time, some experiments utilize Llama-3.0 models instead. The final models described in Section 4 are based on Llama-3.1.
3.1 Learning Rate
Maybe the most important hyperparameter to consider is the maximum learning rate $LR_{max}$ of the continued pretraining. meta have used a $LR_{max}$ of 3.0e-4 and 1.5e-4 for their training of Llama-3.1 8B and 70B respectively Dubey et al. (2024). However, using the same maximum learning rate for continued pretraining might not yield the best results as the model might forget too much information from the previous training or contrary not learn enough from the new data mixture Gupta et al. (2023); Ibrahim et al. (2024).
There are mainly 2 paradigms in existing work: (i) use the same $LR_{max}$ as for the original pretraining Rozière et al. (2023); Chen et al. (2023); Shao et al. (2024), or (ii) use a smaller value, typically around 10% of the original $LR_{max}$ Azerbayev et al. (2024); Lewkowycz et al. (2022); Colombo et al. (2024a); Yuan et al. (2024); Thulke et al. (2024); Xie et al. (2024); Parmar et al. (2024).
$LR_{max}$ ppl ($\downarrow$) benchmark ($\uparrow$)
e-com. general e-com. general
Llama-3.0 7.28 8.38 45.9 44.1
3.0e-5 2.03 6.43 59.9 42.0
3.0e-4 2.01 6.48 58.6 40.5
3.0e-3 2.15 7.70 50.6 34.5
Table 2: Effect of the maximum learning rate of the continued pretraining (1 trillion tokens) of Llama-3.0 8B on the final model performance. Llama-3.0 used $LR_{max}$=3.0e-4.
We perform a set of experiments to determine the best maximum learning rate. Since the impact of the learning rate might significantly depend on the amount of data used in training, we decide to compare training runs utilizing the full 1 trillion tokens of data (50% e-commerce ratio). In all cases, the learning rate decays over the course of the full training with a cosine scheduling to the minimum learning rate of 3.0e-6. We compare the final model performance in terms of perplexity on the heldout test sets, as well as general (average of NLU En and LLM Leaderboard) and e-commerce benchmarks. The results can be found in Table 2.
In terms of perplexity, we find that a higher learning rate leads to a slightly better score on the new domain. However, these improvements do not translate to a better score on the e-commerce specific benchmarks. At the same time, a higher learning rate leads to more degradation on the general domain benchmarks. This might be an indication that our general domain data mix is maybe a bit lower quality than what has been used by meta in the Llama-3 pretraining. In the end, we decide to use an $LR_{max}$ that is 10% of the maximum learning rate used in pretraining, i.e. 3.0e-5 for the 8B model and 1.5e-5 for the 70B model.
3.2 Data Weighting
While we want the model to learn about the new domain, at the same time we want to avoid the effect of catastrophic forgetting. To combat this, it is common to include some percentage of general domain examples in the data mixture (sometimes called ‘replay examples’). Most existing works use only up to 15% of general data in their mixture Azerbayev et al. (2024); Lewkowycz et al. (2022); Rozière et al. (2023); Chen et al. (2023); Colombo et al. (2024b, a) with the exception of Yuan et al. (2024) who use 35%. However, we have reason to believe that in our specific case, a higher ratio of general domain data might be advisable, because our in-domain e-commerce data comes from a very different distribution.
Following Ibrahim et al. (2024) we perform experiments on a limited amount of training tokens (ca 30 billion) with a varying ratio of e-commerce data ($LR_{max}$=3.0e-5). The results can be found in Table 3.
% e-com ppl ($\downarrow$) benchmark ($\uparrow$)
e-com. general e-com. general
Llama-3.0 7.28 8.38 45.9 44.1
10 2.75 6.87 55.6 43.2
25 2.59 6.92 56.7 43.1
50 2.47 7.00 57.5 43.3
75 2.40 7.15 57.6 43.2
Table 3: Effect of the amount of e-commerce data in the continued pretraining (30 billion tokens) of Llama-3.0 8B on the final model performance.
As expected, we see that with a higher percentage of e-commerce data, the perplexity and downstream performance for the e-commerce related tasks is improving, although with diminishing returns when going above 50%. At the same time, perplexity on the general domain data is getting worse, but this does not effect the model scores on the general domain benchmarks. As mentioned before, the reason for this is most likely the different distributions of the general domain data we are using vs the one that was used in pretraining. In the end, we decide to continue pretraining with an e-commerce percentage of 50% in our data mix.
3.3 Context Size
Figure 1: Perplexity as a function of the input sequence length for the 8B (e)-Llama-3.0/3.1 models. The 3.0 variants can not handle context sizes much longer than 8k, since they have never seen these lengths in training.
Figure 2: Evolution of the model performance over the course of the training. Left: perplexity on heldout data sets in the general/e-commerce domain. Center/Right: Evolution of downstream model performance for 8B/70B model.
Finally, we explore the effect of the continued pretraining on the context size of the model. While Llama-3.0 has a context size of 8k, the Llama-3.1 models have a much larger context size of 128k. Ideally we would like to continue the pretraining with the same large context size, but this introduces several challenges. First, the vast majority of our training examples both for general and e-commerce domain are shorter than 1k tokens. Additionally, increasing the context size makes it harder to train the model efficiently due to the quadratic computational complexity of the transformer model. There exist methods to mitigate the latter issue, like the context parallel training approach Fang and Zhao (2024) but when applying said approach, we find that this still introduces too much computational overhead and significantly slows down the training. We decide to continuously train both Llama-3.0 8B and 3.1 8B with 8k context size and study the effects this has on the models. In Figure 1 we calculate the perplexity of the 8B base and e-Llama models as a function of the input sequence length for a general domain heldout test set.
All models exhibit nearly identical performance for inputs smaller than 8k. Unsurprisingly, the 3.0 variants can not handle inputs larger than 8k at all. The e-Llama model that is based on Llama-3.1 exhibits a much better understanding of longer sequences, even though it has not seen any sequences longer than 8k in the continued pretraining. We can conclude that the model retains most of its ability to handle longer sequences. We do see some degradation for even longer input lengths, but this is an acceptable trade-off for us. We therefore decide to perform the continued pretraining with a context size of 8k.
4 e-Llama
In this section we discuss the training and performance of the final e-Llama 8B and 70B models.
4.1 Training
Our setup mostly follows Herold et al. (2024) while taking into account our findings from Section 3. In particular we use cosine Learning Rate (LR) scheduling with warmup, a batch-size of ca. 11.8 million tokens and 85k total update steps.
In Figure 2, we show the evolution of the 8B/70B model performance over the course of the training. We see that the perplexity on the general domain data is decreasing for both 8B and 70B model. At the same time, the gap between 8B and 70B stays constant throughout the training. This indicates that while the distribution of our general domain data is different from the original one, the complexity of the data might be similar. Perplexity on the e-commerce data is also decreasing but at a much faster rate. In the end, the difference in terms of e-commerce perplexity for 8B and 70B model is much smaller than for the base models.
In terms of downstream performance, we find that the 8B model seems to quickly become saturated and performance is no longer increasing after 20% of the training. The 70B model on the other hand recovers much better on the general domain tasks, while also continuously improving in the e-commerce domain. We think this might be due to the much larger model size, that allows the model to better incorporate new information without catastrophic forgetting. Also, the smaller learning rate for the 70B model might have played a role here.
4.2 Final models
In Table 4 we show the final model performance of e-Llama 8B/70B in comparison to the Llama-3.1 base models.
Model general domain benchmarks ($\uparrow$) e-commerce benchmarks ($\uparrow$)
En non-En En non-En
[…]
8B
Llama-3.1 71.8 17.2 54.1 36.5 61.8 50.1 27.4 55.3 35.8
e-Llama 71.6 12.6 54.0 54.9 74.9 59.6 37.8 67.4 46.8
70B
Llama-3.1 76.6 29.7 58.5 42.8 66.3 59.3 35.2 61.9 40.4
e-Llama 76.3 28.7 59.2 59.2 79.5 65.7 49.9 71.5 52.8
Table 4: Final performance of the e-Llama 8B/70B models on general domain and e-commerce specific evaluation benchmarks.
On the general domain (non-En) NLU benchmarks, the e-Llama models perform the same as the base Llama-3.1 models. On the more challenging LLM Leaderboard tasks, we see some performance degradation, especially for the smaller 8B model variant. We think this might be due to a combination of smaller model size and a different data distribution of our general domain data compared to what has been used at meta.
On the e-commerce benchmarks, the e-Llama models improve relative to the Llama-3.1 base models by around 25% on English and by around 30% on non-English benchmarks on average. Interestingly, the gap between 8B and 70B variant for Llama base model and for e-Llama is roughly the same for the e-commerce tasks, even though in terms of perplexity the e-Llama models are closer together (compare left side of Figure 2). This once again highlights that perplexity and downstream performance do not always follow the same trend.
4.3 Model Merging
The last topic we want to discuss is how to better align the trade-off between general and domain-specific performance. Lets assume we want to have the best performing model on the e-commerce domain, but we can not allow the general domain performance to drop below a certain threshold on the general domain tasks. As can be seen from the experiments in Figure 2 and Table 3, reducing the percentage or amount of e-commerce training data does not allow to make very precise forecasts of final model performance. Instead, we utilize a technique called ‘model merging’ Wortsman et al. (2022), where we simply average all parameters of the base Llama-3.1 model checkpoint and our final e-Llama model checkpoint. In Figure 3 we show how the performance of the resulting model changes as a function of the individual model weights.
Figure 3: Model merging: 8B Model performance on general and e-commerce benchmarks as a function of the weight of the e-Llama model parameters vs the base Llama-3.1 model parameters.
The performance for both general and e-commerce domain follow an almost linear trend. This allows for a very precise tuning of the final model performance and has the additional advantage that the model merging is not compute intensive at all.
5 Conclusion
We have discussed our efforts to adapt the Llama-3.1 8B and 70B parameter base models towards the e-commerce domain. In order to evaluate the model capabilities in the e-commerce setting, we design and implement a set of multilingual, e-commerce specific evaluation benchmarks. Through a series of experiments, we determine the best experimental setting for our use-case. We show that the models can be adapted well towards the new domain with limited degradation on general domain performance. Furthermore, we highlight that with model merging, we can very precisely tune the final model performance.
Appendix A Appendix
A.1 Examples for e-commerce tasks
Here, we give an example for each of the e-commerce tasks described in Section 2.3. All examples are for the English language. For languages other than English, the prompt stays the same, but the item-specific attributes like title are in the corresponding language.
AP
The model has to predict the most probable continuation of the following text input:
For an e-commerce website, under the category "Video Games & Consoles:Video Games", the listing with the title "Dark Souls III (Sony PlayStation 4)" has the following aspect key-value pairs:
Rating:
[…]
The model is used to independently score the following 4 text sequences, has to give the highest probability to the correct sequence (first one).
For an e-commerce website, the listing with the title "Dark Souls III (Sony PlayStation 4)" has the following aspect key-value pairs associated with it:
Rating: M - Mature
For an e-commerce website, the listing with the title "Dark Souls III (Sony PlayStation 4)" has the following aspect key-value pairs associated with it:
Rating: E - Everyone
For an e-commerce website, the listing with the title "Dark Souls III (Sony PlayStation 4)" has the following aspect key-value pairs associated with it:
Rating: T - Teen
For an e-commerce website, the listing with the title "Dark Souls III (Sony PlayStation 4)" has the following aspect key-value pairs associated with it:
Rating: AO - Adults Only
[…]
The model is used to independently score the following 4 text sequences, has to give the highest probability to the correct sequence (first one).
For the listing with the title "Authentic Louis Vuitton Monogram Empreinte Bastille PM 2Way Tote Bag Black 9281E", the final selling price was \$816.00.
For the listing with the title "Authentic Louis Vuitton Monogram Empreinte Bastille PM 2Way Tote Bag Black 9281E", the final selling price was \$81.60.
For the listing with the title "Authentic Louis Vuitton Monogram Empreinte Bastille PM 2Way Tote Bag Black 9281E", the final selling price was \$204.00.
For the listing with the title "Authentic Louis Vuitton Monogram Empreinte Bastille PM 2Way Tote Bag Black 9281E", the final selling price was \$1632.00.
MCA
The model has to predict the most probable continuation of the following text input:
For an e-commerce website, under the category "Clothing, Shoes & Accessories:Women:Women’s Clothing:Coats, Jackets & Vests", the following are the most common aspect values for the aspect key "Outer Shell Material":
${}^{\text{MC}}$
The model is used to independently score the following 4 text sequences, has to give the highest probability to the correct sequence (first one).
For an e-commerce website, under the category "Cameras & Photo:Digital Cameras", the most common aspect value for the aspect key "Brand" is "Canon".
For an e-commerce website, under the category "Cameras & Photo:Digital Cameras", the most common aspect value for the aspect key "Brand" is "Fujifilm".
For an e-commerce website, under the category "Cameras & Photo:Digital Cameras", the most common aspect value for the aspect key "Brand" is "PENTAX".
For an e-commerce website, under the category "Cameras & Photo:Digital Cameras", the most common aspect value for the aspect key "Brand" is "Nikon".
A.2 Optimizing the Few-Shot Setup
For the base Llama-3 8B model, we prompt the model for each of the above tasks with up to 20 few-shot examples. The results can be seen in Figure 4.
Figure 4: Llama-3 8B model performance on the 5 e-commerce evaluation tasks as a function of the number of few-shot examples provided in the prompt.
[…]